# Broadway: Boost Your Text-to-Video Generation Model in a Training-free way

### v1

@@ A.1.1 Frequency Manipulation
Given a temporal attention map $\mathcal{A}\in\mathbb{R}^{(B\times H\times W)\times F\times F}$ with batch size $B$, spatial resolution $H\times W$ and frame number $F$, since we treat it as a batch of attention sequences, we will next discuss the operations performed on a single softmax sequence $x[n]$ of length $F$.
@@
$X[k]=\sum_{n=0}^{F-1}x[n]\cdot e^{-j\frac{2\pi}{N}kn},~{}k=0,1,\dots,F-1,$ (8)
@@
$E_{x}=\sum_{n=0}^{F-1}x[n]^{2}=\frac{1}{F}\sum_{k=0}^{F-1}X[k]^{2},$ (9)
As mentioned in Section 4.2.2, Fourier-based Motion Enhancement uses a threshold index $\tau$ to separate the high-frequency and low-frequency components of the sequence, scaling the high-frequency components by a factor of $\beta$. This operation can be expressed as:
@@
$=\frac{(\beta^{2}-1)}{F}\sum_{k\in[\frac{F}{2}-\tau,\frac{F}{2}+\tau]}X^{2}[k],$
@@
Furthermore, it can be demonstrated that the attention sequence processed by Fourier-based Motion Enhancement remains a softmax sequence. This property is preserved because the DC component $X[0]$ of the attention sequence, which determines the sum of the sequence, is not modified throughout the operation. By plugging $k=0$ into Eq. 8, we can ascertain this property:
$X[0]=\sum_{n=0}^{F-1}x[n]=\sum_{n=0}^{F-1}x^{{}^{\prime}}[n]=1,$ (12)
A.1.2 adaptive $\beta$
As depicted in Fig. 7, let $E_{1}$ denote the the energy of the temporal attention map before applying BroadWay operations, $E_{2}$ the energy after Temporal Self-Guidance, and $E_{3}$ the energy after Fourier-based Motion Enhancement. Here, we demonstrate that using the adaptive $\beta$ as defined in Eq. 7 ensures that $E_{3}\geq E_{1}$.
Based on the separation of high-frequency and low-frequency components in the sequence as described in Section A.1.1, we can compute the energy of the high-frequency and low-frequency parts of the sequence $x[n]$, denoted as $E_{x}^{H}$ and $E_{x}^{L}$, respectively:
@@
$E_{x}^{H}$ $=\frac{1}{F}\sum_{k\notin[\frac{F}{2}-\tau,\frac{F}{2}+\tau]}X^{2}[k],$
@@
$E_{x}=E_{x}^{H}+E_{x}^{L},$ (14)
Furthermore, we can concisely express the energy manipulation performed by Fourier-based Motion Enhancement described in Section A.1.1, as follows:
@@
$\beta_{c}=\sqrt{\frac{E_{1}-E_{2}^{L}}{E_{2}^{H}}},$ (18)
In BroadWay operations, the user-specified $\beta$, denoted as $\beta_{0}$, will be compared with the critical value $\beta_{c}$, and the larger of the two will be selected as the actual $\beta$ value in Fourier-based Motion Enhancement:
$\beta=\left\{\begin{array}[]{rcl}\beta_{0}&&{\beta_{0}\geq\beta_{c}},\\ \beta_{c}&&{\beta_{0}<\beta_{c}},\end{array}\right.$ (19)
By adopting such a adaptive $\beta$ value, it can be theoretically guaranteed that the energy of the temporal attention map is increased during BroadWay  operations, thereby enhancing the motion magnitude in synthesized videos.
A.2 Additional Qualitative Results
In this section, we provide additional qualitative comparison results of BroadWay on AnimateDiff (Fig. 12, Fig. 13, Fig. 17 and Fig. 14) and VideoCrafter2 (Fig. 15, Fig. 16, Fig. 18).
A.3 Materials used in quantitative experiments
User Study Details. In our user study, each participant receives 50 videos synthesized by Vanilla T2V backbones and 50 videos synthesized by BroadWay-enhanced backbones. These videos are sampled from the same random seeds to ensure fair comparison. For each video pair from Vanilla and Vanilla+BroadWay, participants are required to select the video they perceive as superior based on overall Video Quality, considering both structure coherence and motion magnitude, and cast their vote accordingly. The videos were presented in a randomized order to reduce potential bias, and participants were allowed ample time to review each pair before making their selections.
MLLM Prompt. Here, we present the prompt used in the MLLM assessment.
    query = """
    You are provided with two sets of video frames, each containing
    4 representative frames, along with a shared textual prompt that
    was used to generate both videos.
    Your task is to perform a comparative evaluation of the two videos,
    focusing on their structure rationality / motion consistency.
    """.strip()
    prefix_1 = """
    Here is the frame data of video_1:
    prefix_2 = """
    Here is the frame data of video_2:
    suffix = """
    Based on your evaluation of motion consistency, choose the video
    set you find to be superior.
    If you determine that the first set of frames (Video_1) is better,
    respond with "A". If the second set (Video_2) is superior, respond
    with "B". Return only "A" or "B" based on your assessment.
A.4 Limitations
Parameter Sensitivity. The default values of BroadWay parameters $\alpha$ and $\beta$ are relatively robust within a specific T2V backbone but may not be universally optimal for different backbones. Users seeking enhanced visual quality are encouraged to manually adjust these parameters. Increasing $\alpha$ can lead to stronger motion dynamics, while a higher value of $\beta$ enhances structural consistency.
Performance Upper Bound. Although BroadWay demonstrates the capability to unlock the synthesis potential of various T2V backbones, the synthesized videos remain confined within the sampling distribution of the original T2V backbone. Therefore, the upper performance bound of our proposed method is still constrained by the original T2V backbone.
Figure 12: More results on AnimateDiff (Object Motion Enhancement).
Figure 13: More results on AnimateDiff (Object Motion Enhancement).
Figure 14: More results on AnimateDiff (Corrupted Case Repair).
Figure 15: More results on VideoCrafter2 (Object Motion Enhancement).
Figure 16: More results on VideoCrafter2 (Corrupted Case Repair).
Figure 17: More results on AnimateDiff (Camera Motion Enhancement).
Figure 18: More results on VideoCrafter2 (Camera Motion Enhancement).

### v3

@@ 10.1 Frequency Components Manipulation
Given a temporal attention map $\mathcal{A}\in\mathbb{R}^{(B\times H\times W)\times F\times F}$ with batch size $B$, spatial resolution $H\times W$ and frame number $F$, since we treat it as a batch of 1D attention sequences, we will next discuss the operations performed on a single softmax sequence $x[n]$ of length $F$.
@@
$X[k]=\sum_{n=0}^{F-1}x[n]\cdot e^{-j\frac{2\pi}{N}kn},~{}k=0,1,\dots,F-1.$ (8)
@@
$E_{x}=\sum_{n=0}^{F-1}x[n]^{2}=\frac{1}{F}\sum_{k=0}^{F-1}X[k]^{2}.$ (9)
As mentioned in the main paper, Fourier-based Motion Enhancement uses a threshold index $\tau$ to separate the high-frequency and low-frequency components of the sequence, scaling the high-frequency components by a factor of $\beta$. This operation can be expressed as:
@@
$=\frac{(\beta^{2}-1)}{F}\sum_{k\in[\frac{F}{2}-\tau,\frac{F}{2}+\tau]}X^{2}[k].$
@@
Furthermore, it can be demonstrated that the attention sequence processed by Fourier-based Motion Enhancement remains a softmax sequence. This property is preserved because the direct current (DC) component $X[0]$ of the attention sequence, which determines the sum of the sequence, is not modified throughout the operation. By plugging $k=0$ into Eq. 8, we can ascertain this property:
$X[0]=\sum_{n=0}^{F-1}x[n]=\sum_{n=0}^{F-1}x^{{}^{\prime}}[n]=1.$ (12)
10.2 Adaptive $\beta$ in ByTheWay Operations
Figure 12: ByTheWay Operations.
As depicted in the Fig. 12, let $E_{1}$ denote the the energy of the temporal attention map before applying ByTheWay operations, $E_{2}$ the energy after Temporal Self-Guidance, and $E_{3}$ the energy after Fourier-based Motion Enhancement. Here, we demonstrate that using the adaptive $\beta$ as defined in Eq. 13 ensures that $E_{3}\geq E_{1}$.
$\beta(E_{1},E_{2})=max\{\beta_{0},\sqrt{\frac{E_{1}-E_{2}^{L}}{E_{2}^{H}}}\},$ (13)
Based on the separation of high-frequency and low-frequency components in the sequence as described in Section 10.1, we can compute the energy of the high-frequency and low-frequency parts of the sequence $x[n]$, denoted as $E_{x}^{H}$ and $E_{x}^{L}$, respectively:
@@ 10.2 Adaptive $\beta$ in ByTheWay Operations
$E_{x}^{H}$ $=\frac{1}{F}\sum_{k\notin[\frac{F}{2}-\tau,\frac{F}{2}+\tau]}X^{2}[k].$
@@
$E_{x}=E_{x}^{H}+E_{x}^{L}.$ (15)
Furthermore, we can concisely express the energy manipulation performed by Fourier-based Motion Enhancement described in Section 10.1, as follows:
@@
$\beta_{c}=\sqrt{\frac{E_{1}-E_{2}^{L}}{E_{2}^{H}}}.$ (19)
In ByTheWay operations, the user-specified $\beta$, denoted as $\beta_{0}$, will be compared with the critical value $\beta_{c}$, and the larger of the two will be selected as the actual $\beta$ value in Fourier-based Motion Enhancement:
$\beta=max\{\beta_{0},\beta_{c}\}.$ (20)
By adopting such a adaptive $\beta$ value, it can be theoretically guaranteed that the energy of the temporal attention map is increased during ByTheWay operations, thereby enhancing the motion magnitude in synthesized videos.
11 Limitation
Although ByTheWay demonstrates the capability to unlock the synthesis potential of various T2V backbones, the synthesized videos remain confined within the sampling distribution of the original T2V backbone. Therefore, one limitation of our method is that its performance upper bound is still constrained by the original T2V backbone.
Figure 13: Results on CogVideoX.
Figure 14: Ablation on Guidance Anchor.
Figure 15: Ablation on More Sampling Steps.
Figure 16: Ablation on Operation Steps. Prompt: “a vintage car drives on a country road, …”
Figure 17: More Results on AnimateDiff (Motion Enhancement).
Figure 18: More Results on AnimateDiff (Structure Enhancement).
Figure 19: More Results on VideoCrafter2 (Motion Enhancement).
Figure 20: More Results on VideoCrafter2 (Structure Enhancement).